OpenStudy (anonymous):

Any idea on how to do this problem: √27x^2y-√45xy^2+√108x^2y-√125xy^2

OpenStudy (anonymous):

\[\sqrt{27x^2y}-\sqrt{45xy^2}+\sqrt{108x^2y}-\sqrt{125xy^2}\] \[=\sqrt{3^23x^2y}-\sqrt{3^23xy^2}+\sqrt{6^23x^2y}-\sqrt{5^25xy^2}\] Simplify taking out the squares: \[=3x\sqrt{3y}-3y\sqrt{3x}+6x\sqrt{3y}-5y\sqrt{5x}\] Combine like types: \[=9x\sqrt{3y}-3y\sqrt{3x}-5y\sqrt{5x}\] You can also rewrite this as: \[=9x\sqrt{3y}-y(3\sqrt{3x}+5\sqrt{5x})\]

OpenStudy (anonymous):

9x√3y-8y√5x